OpenStudy (anonymous):

Derivative of 8-x^3, Step by step please.

OpenStudy (lgbasallote):

use the power rule. x^n = nx^n-1 and the constant rule which is derivative of k (constant) = 0 so you have a constant and a power apply the constant rule to 8, it will become zero apply the power rule to -x^3, it will become -3x^(3-1) so the derivative of -x^3 = -3x^2 combine them 0 -3x^2 the derivative = -3x^2